OpenStudy (anonymous):

What is the distance from (−2, 4) to (0, −6)? Round your answer to the nearest hundredth. 9.23 10.20 10.98 11.34

OpenStudy (cookielate):

d=(x2−x1)2+(y2−y1)2−−−−−−−−−−−−−−−−−−√ d=(0+2)2+(−6−4)2−−−−−−−−−−−−−−−−√ Add & Subtract: d=(2)2+(−10)2−−−−−−−−−−−√ Simplify Exponents: d=4+100−−−−−−√ Add: d=104−−−√ Can you square that

OpenStudy (cookielate):

if its too hard the answer is B